# Geometry of convex geometries

### v1

@@ 5.3 Realization of ideals of convex geometries and VC-dimension
If $\mathcal{A}=(U,\mathcal{C})$ is an ideal semialignment with $k$ facets or a median set systems with $k$ maxima, then $\mathcal{A}$ is realizable and $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})\leq\min(|U|,k\operatorname{dim_{VC}}(\mathcal{A}))$.
@@
From Corollaries 23 and 24 and Theorem 25 we get realizability and the dimension bound of $|U|$. Now, using Corollary 37 we obtain the bounds of the form $k\operatorname{dim_{VC}}(\mathcal{A}))$. ∎
In this subsection we confirm 38 for trees and multisimplicial complexes (alias, ideals of the grid). Indeed, both our bounds are best-possible.
Trees are the simplest examples of ideals of convex geometries that are not convex geometries. They are also the median set systems of VC-dimension 1.
@@
Median set systems are realizable by Corollary 39. Since median systems are ample, their VC-dimension coincides with the dimension of a largest cube, and thus with their (topological) dimension. Corollary 37 and Proposition 40 motivate the following strengthening of 38 in the case of median systems:
@@
Now, suppose that some cell $C$ of $\mathcal{H}_{1}$ represents an ideal of $X$ that does not correspond to an element of $L_{1}$. Since $C$ and $P$ are open and convex we can choose a hyperplane $H$ that separates $C$ and $P$ and intersects the closure $\overline{P}$ of $P$ in $\overline{P}\cap\overline{C}$. Thus, if we add to $K$ the halfspace defined by $H$ that contains $P$, then we remove $C$ from the arrangement and all cells of $\mathcal{H}_{1}$ that correspond to faces not included in $\overline{P}\cap\overline{C}$ are still present. The only further cells of $\mathcal{H}_{1}$ that are removed are those that correspond to ideals of $X$ corresponding to supersets of the ideal corresponding do $C$. Since $L_{1}$ is an ideal, we did not want these regions. Consequently, taking as $K$ the intersection of all open halfspaces containing $P$ and defined by hyperplanes separating $P$ from cells $C$ representing ideals of $X$ that does not correspond to elements of $L_{1}$, we will obtain that $L_{1}=\mathcal{M}(\mathcal{H}_{1},K)$. ∎
Case 44.
@@
By induction hypothesis, we can define the dilates $P_{1},P_{2},\ldots,P_{i-1}$ of $P$, their arrangements of supporting hyperplanes $\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{i-1}$, and the convex set $K$ in such a way that $\mathcal{M}(\bigcup_{j=1}^{i-1}\mathcal{H}_{j},K)=L_{i-1}$. First, suppose that there exists an element $x\in L_{i}$ that is represented by a cell $C$ touching $P_{i}$ and that lies outside of $K$. Since $L$ is an ideal, the element $y\in L_{i-1}$ cutting the ideal of $x$ to height at most $i-1$ on each chain is in $L$. In particular, this point $p$ is represented by a cell touching $P_{i-1}$ within $K$. Hence, if $q$ is a point in the cell representing $x$, a line from $p$ to $q$ must cross the boundary of $K$. We can thus dilate $P_{i}$ to a smaller polytope such that the cell representing $x$ intersects the interior of $K$. In order to avoid now some of the regions touching $P_{i}$ from the outside, we proceed as in the case of $P_{1}$ noting that all the hyperplanes that we add to $K$ are disjoint from the interior of $P_{i}$ hence they do not interfere with the so far obtained representation of $L_{i-1}$. ∎
@@ 6 Closing remarks
6.1 Concepts of dimension
We have shown that the Euclidean dimension $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ of a convex geometry $\mathcal{A}=(U,\mathcal{C})$ lies between its VC-dimension $\operatorname{dim_{VC}}(\mathcal{A})$ and the minimum of $|U|$ and $\operatorname{cdim}(\mathcal{A})$. It is known that convex geometries with bounded convex dimension and unbounded $|U|$ exist. On the other hand also there are convex geometries with $\operatorname{cdim}(\mathcal{A})$ exponential in $|U|$, see [19]. Hence none of our bounds on $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ dominates the other. It would be interesting to compare $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ with other concepts of dimension for convex geometries, e.g., Dushnik-Miller dimension, Boolean dimension, local dimension, and fractional dimension. Their behavior on convex geometries has been studied recently in [19].
Figure 3: A bouquet of convex geometries that is not an ideal of a convex geometry. The bouquets of convex geometries is represented in black as isometric subgraph of the 4-dimensional hypercube (in gray)
6.2 From ideals to bouquets
To finish we want to extend our vision from ideals of convex geometries. A pair $\mathcal{A}=(U,\mathcal{C})$ of a finite universe $U$ and a collection of convex sets $\mathcal{C}\subseteq 2^{U}$ is called bouquet of convex geometries if it satisfies the following two conditions:
  • (C1)
    $\varnothing\in\mathcal{C}$;
  • (C2)
    $X\cap Y\in\mathcal{C}$ for all $X,Y\in\mathcal{C}$;
  • (C3)
    for all $X,Y\in\mathcal{C}$ with $Y\subset X$, there is $e\in X\setminus Y$ such that $Y\cup e\in\mathcal{C}$; (locally extendable)
These structures were called conditional antimatroids in [2]. It is easy to see that ideals of convex geometries are bouquets of convex geometries. Note that in a sense ideal semialignments were defined as bouquets of convex geometries with the additional property of being locally union-closed. We then showed that they are indeed ideals of convex geometries in Proposition 22. However, there are bouquets of convex geometries that are not ideals.
Proposition 46.
Proof.
Consider the bouquet of convex geometries in Figure 3, right, i.e., $E^{\prime}=\{1,2,3,4\}$ and
Suppose that $\mathcal{A}^{\prime}=(E^{\prime},\mathcal{C}^{\prime})$ is an ideal of a convex geometry $\mathcal{A}=(E,\mathcal{C})$. One can remove elements from $E^{\prime}$ until $E=E^{\prime}$. Hence, $E\in\mathcal{C}^{\prime}$ by (C1). By (C3) applied to $\{1,4\}$, one of the sets $\{1,2,4\}$ or $\{1,3,4\}$ must be in $\mathcal{C}$. Now applying (C2) with respect to $\{1,2,4\}$ and $\{2,3,4\}$ or $\{1,3,4\}$ and $\{1,2,3\}$, we get that $\{2,4\}$ or $\{1,3\}$ are in $\mathcal{C}$. Hence $\mathcal{A}^{\prime}=(E^{\prime},\mathcal{C}^{\prime})$ is not an ideal of $\mathcal{A}$. ∎

### v2

@@ 5.3 Realization of ideals of convex geometries and VC-dimension
If $\mathcal{A}=(U,\mathcal{C})$ is a bouquet of downset alignments with $\ell$ maxima, e.g., a median set systems with $\ell$ maxima, then $\mathcal{A}$ is realizable and $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})\leq\min(|U|,\ell\operatorname{dim_{VC}}(\mathcal{A}))$.
@@
From Theorem 23, Corollary 25, and Theorem 26 we get realizability and the dimension bound of $|U|$. Now, using Theorem 23 and Corollary 38 we obtain the bounds of the form $\ell\operatorname{dim_{VC}}(\mathcal{A}))$. ∎
In this subsection we confirm 39 for trees and multisimplicial complexes (alias, ideals of the grid). Indeed, both our bounds are best-possible.
A tree is a connected acyclic graph. A rooted tree is a tree with a distinguished vertex $r$. A family of sets $\mathcal{A}=(U,\mathcal{C})$ is a tree if $\mathcal{A}$ is a median system whose 1-inclusion graph is a tree.
@@
Median set systems are realizable by Corollary 40. Since median systems are ample, their VC-dimension coincides with the dimension of a largest cube, and thus with their (topological) dimension. Corollary 38 and Proposition 41 motivate the following strengthening of 39 in the case of median systems:
@@
Now, suppose that some cell $C$ of $\mathcal{H}_{1}$ represents an ideal of $X$ that does not correspond to an element of $L_{1}$. Since $C$ and $P$ are open and convex we can choose a hyperplane $H$ that separates $C$ and $P$ and intersects the closure $\overline{P}$ of $P$ in $\overline{P}\cap\overline{C}$. Thus, if we add to $K$ the halfspace defined by $H$ that contains $P$, then we remove $C$ from the arrangement and all cells of $\mathcal{H}_{1}$ that correspond to faces not included in $\overline{P}\cap\overline{C}$ are still present. The only further cells of $\mathcal{H}_{1}$ that are removed are those that correspond to ideals of $X$ corresponding to supersets of the ideal corresponding do $C$. Since $L_{1}$ is an ideal, we did not want these regions. Consequently, taking as $K$ the intersection of all open halfspaces containing $P$ and defined by hyperplanes separating $P$ from cells $C$ representing ideals of $X$ that does not correspond to elements of $L_{1}$, we will obtain that $L_{1}=\mathcal{M}(\mathcal{H}_{1},K)$. $\Diamond$
Case 2.
@@
By induction hypothesis, we can define the dilates $P_{1},P_{2},\ldots,P_{i-1}$ of $P$, their arrangements of supporting hyperplanes $\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{i-1}$, and the convex set $K$ in such a way that $\mathcal{M}(\bigcup_{j=1}^{i-1}\mathcal{H}_{j},K)=L_{i-1}$. First, suppose that there exists an element $x\in L_{i}$ that is represented by a cell $C$ touching $P_{i}$ and that lies outside of $K$. Since $L$ is an ideal, the element $y\in L_{i-1}$ cutting the ideal of $x$ to height at most $i-1$ on each chain is in $L$. In particular, this point $p$ is represented by a cell touching $P_{i-1}$ within $K$. Hence, if $q$ is a point in the cell representing $x$, a line from $p$ to $q$ must cross the boundary of $K$. We can thus dilate $P_{i}$ to a smaller polytope such that the cell representing $x$ intersects the interior of $K$. In order to avoid now some of the regions touching $P_{i}$ from the outside, we proceed as in the case of $P_{1}$ noting that all the hyperplanes that we add to $K$ are disjoint from the interior of $P_{i}$ hence they do not interfere with the so far obtained representation of $L_{i-1}$. $\Diamond$
@@ 6 Closing remarks
Concepts of dimension.
We have shown that the Euclidean dimension $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ of a convex geometry $\mathcal{A}=(U,\mathcal{C})$ lies between its VC-dimension $\operatorname{dim_{VC}}(\mathcal{A})$ and the minimum of $|U|$ and $\operatorname{cdim}(\mathcal{A})$. It is known that convex geometries with bounded convex dimension and unbounded $|U|$ exist. On the other hand also there are convex geometries with $\operatorname{cdim}(\mathcal{A})$ exponential in $|U|$, see [22]. Hence none of our bounds on $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ dominates the other. We believe that $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ can be bounded by a function of $\operatorname{dim_{VC}}(\mathcal{A})$, see 39. Further, it would be interesting to compare $\operatorname{dim_{\mathbb{E}}}(\mathcal{A})$ with other concepts of dimension for convex geometries, e.g., Dushnik-Miller dimension, Boolean dimension, local dimension, and fractional dimension. Their behavior on convex geometries has been studied recently in [22]. Another direction are representations by spheres and ellipsoids as studied in [1]
From ideals to bouquets.
@@
Example 46.
The bouquet of convex geometries $(U,\mathcal{C}^{\prime})$ of the right of Figure 1 is realizable. The difference between $(U,\mathcal{C}^{\prime})$ and the convex geometry $(U,\mathcal{C})$ of the left of Figure 1 is that $\{1,2,3,4\}\in\mathcal{C}\setminus\mathcal{C}^{\prime}$ and $\{1,4\}\in\mathcal{C}^{\prime}\setminus\mathcal{C}$. Therefore, one cannot obtain a representation of $(U,\mathcal{C}^{\prime})$ by adding linear constraints to the representation of $(U,\mathcal{C})$.
However, we claim that $(U,\mathcal{C}^{\prime})$ is realized in $\mathbb{R}^{4}$ by the convex set $K$ defined by the following two inequalities:
$x_{1}+3x_{3}$ $<x_{2}$ $3x_{2}+x_{4}$ $<x_{3}$
Observe that there is no point $x$ in the realization such that $x_{1}>0$, $x_{3}>0$, and $x_{2}<0$ (respectively, $x_{2}>0$, $x_{4}>0$, and $x_{3}<0$). Consequently, if $X$ is one of the sets $\{1,3\}$ or $\{1,3,4\}$ (respectively, $\{2,4\}$ or $\{1,2,4\}$), then $K$ does not contain any point of the $X$-orthant. Moreover, by adding the two inequalities, we get $x_{1}+2x_{2}+2x_{3}+x_{4}<0$, and thus for $X=\{1,2,3,4\}$, there is no point $x\in K$ in the $X$-orthant. Therefore, the convex set $K$ does not intersect any $X$-orthant such that $X\notin\mathcal{C}^{\prime}$. We now consider a set $X\in\mathcal{C}^{\prime}$ and we construct a point of the $X$-orthant that belongs to $K$. If $X$ contains at most two elements, let $x_{i}=1$ for each $i\in X$. We can then find negative values for the remaining coordinates so that both equations are satisfied. Consider now the set $X=\{1,2,3\}$ (respectively, $\{2,3,4\}$) and the point $x=(1,5,1,-15)\in\mathbb{R}^{4}$ (respectively, $(-15,1,5,1)\in\mathbb{R}^{4}$). One can check that $x$ corresponds to $X$ and satisfies both inequalities. Consequently, the bouquet of convex geometries $(U,\mathcal{C})$ is realizable.
Acknowledgment.
We are grateful to the referees for a careful reading and useful remarks that helped us to improve the presentation of the paper.